# Avoided-crossings, degeneracies and Berry phases in the spectrum of quantum noise through analytic Bloch-Messiah decomposition

### v2

@@ 2.  For $k=1,2,\ldots,N$:
Finally, we remark that one can easily adapt Algorithm 2 to obtain a reduced joint-minimum-variation BMD, where only a selected portion of dominant singular values and corresponding singular vectors is computed. This feature lowers the overall computational complexity of the algorithm and has been used in the experiments reported in Section IV.2.
@@
IV Physical examples
In this final section, we apply the theoretical results and algorithms from Section III to two physical systems. We begin with a small system that is small enough to be fully described while still exhibiting all the features of generic systems. We then conclude with a higher-dimensional one.
IV.1 A four-mode system
@@
Figure 1: Case $n=4$ and $\text{codim}=3$. Squeezing, $d_{n+j}^{-2}(\omega)$ and anti-squeezing, $d_{i}^{2}(\omega)$ (with $j\in\{1,\ldots,4\}$) spectra, in dB. The specific choice for the parameters gives an avoided-crossing between the second ($j=1$) and third ($j=2$) squeezing and anti-squeezing spectra around $\omega=0.794\gamma$ (marked by the vertical black-dashed line) and an avoided-crossing between the third ($j=2$) and forth ($j=3$) squeezing spectra around $\omega=0.164\gamma$ (marked by the vertical black-dot-dasjhed line). The inset shows a magnified view of the two avoided crossing regions in the anti-squeezing spectra.
@@
Figure 2: Anti-squeezing spectra in dB ($d_{1}^{2}(\omega,g_{12})$ and $d_{2}^{2}(\omega,g_{12})$) for the case $n=4$, $\text{codim}=3$. (a) For $g_{22}=1.25\gamma$: the surfaces of anti-squeezing avoid crossing through the entire plane $(\omega,g_{11})$. (b) For $g_{22}\approx 1.12\gamma$: the surfaces of squeezing touch at one point of degeneracy in the plane $(\omega,g_{11})$.
@@
Figure 4: Case $n=4$ and $\text{codim}=2$. Squeezing, $d_{n+j}^{-2}(\omega)$ and anti-squeezing, $d_{j}^{2}(\omega)$ (with $j\in\{1,\ldots,4\}$) spectra, in dB. The specific choice for the parameters gives an avoided-crossing between the second ($j=1$) and third ($j=2$) squeezing and anti-squeezing spectra around $\omega=0.761\gamma$ (marked by the vertical black-dashed line) and an avoided-crossing between the third ($j=2$) and forth ($j=3$) squeezing spectra around $\omega=0.079\gamma$ (marked by the vertical black-dot-dashed line). The inset shows a magnified view of the two avoided crossing regions in the anti-squeezing spectra.
@@
We also considered the situation where the couplings are fully real, replacing $F$ of (29) with its real part. Computing the four curves of singular values, we obtained Figure 4, which is qualitatively similar to Figure 1 and shows two avoided crossings. When investigating the existence of a degeneracy near the second avoided crossing as we did previously, we observed that the Berry phases considered in Theorem 3, and computed via Algorithm 2, took only values that are multiples of $\pi$. This behavior contrasts with the codimension 3 case, where, in general, $\alpha_{j}$ varies smoothly across a continuous range of values when $\theta$ is varied (see Figure 3, particularly $\alpha_{j}$ for $j=2,3$). This suggests that the codimension of degeneracies may be lower than expected  [^5] For instance, the Berry phase takes only values that are multiples of $\pi$ when the matrix function is real symmetric, in which case the codimension of degeneracies is in fact 2, see [35]. Indeed, this turns out to be the case. It is easy to see that, when both $F$ and $G$ are real, the matrix $S(\omega)$ satisfies
@@
IV.2 The pulsed microring resonator in synchronously pumped regime
@@
Figure 8: First four most anti-squeezed morphing supermodes ($j={1,\ldots,4}$) for a pulsed microring resonator in the synchronously pumped regime with $n=51$. The enlarged views show the rapid rotation of two supermodes near an avoided crossing.
@@ V Conclusions
Acknowledgements.
The authors thank Prof. Luca Dieci for useful discussions. This work has been partially supported by GNCS–INdAM and by the PRIN2022PNRR n. P2022M7JZW “SAFER MESH - Sustainable mAnagement oF watEr Resources ModEls and numerical MetHods” research grant, funded by the Italian Ministry of Universities and Research (MUR), by the European Union through Next Generation EU, M4.C2.1.1, CUP H53D23008930001 and by Plan France 2030 through the project ANR-22-PETQ-0013 (OQuLus).
@@ Appendix A On the joint-MVD of a complex symmetric matrix-valued function depending on a parameter
where $U,V\in\mathbb{C}^{n\times n}$ are unitary and $\Sigma=\operatorname{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{n})$, with non-negative $\sigma_{j}$’s arranged in non-increasing order. If $A=A^{T}$, $A$ also admits a Takagi decomposition
@@ Appendix B Additional figures on rapid variation of the singular vectors
Here appendix we show the behavior of the morphing supermodes of the case studied in subsection IV.1. In particular, Figure 9 consists of eight panels, each corresponding to a fixed value of $m$, and showing the absolute value of the coefficient $U_{m,j}(\omega)$ across all supermodes $j=\{0,\ldots,7\}$. The vertical black dash-dotted line marks the position of the first avoided crossing at $\omega=0.164\gamma$ occurring between the anti-squeezed supermodes $j=2$ and $j=3$. For the symmetry property of the spectrum of the singular values, the same avoided crossing also occurs between the squeezed supermodes $j=6$ and $j=7$. The vertical black dashed line marks the position of the second avoided crossing at $\omega=0.794\gamma$ occurring between the anti-squeezed supermodes $j=1$ and $j=2$. The same avoided crossing also occurs between the squeezed supermodes $j=5$ and $j=6$. Focusing on the top-left panel of the figure, which displays the absolute value of the $m=0$ coefficient for the eight supermodes, we observe a rapid rotation of the anti-squeezed supermode pair $j=2$ (green line) and $j=3$ (red line) at the first avoided crossing, as well as a rotation of the squeezed pair $j=6$ (pink line) and $j=7$ (gray line). At the second avoided crossing, a similar rotation occurs between the anti-squeezed supermodes $j=1$ (orange line) and $j=2$ (green line), and between the squeezed modes $j=5$ (brown line) and $j=6$ (pink line). The same behavior is observed in the panels corresponding to the other coefficients.

### v3

@@ 2. For $k=1,2,\ldots,N$:
Finally, we remark that one can easily adapt Algorithm 2 to obtain a reduced joint-minimum-variation BMD, where only a selected portion of dominant singular values and corresponding singular vectors is computed. This feature lowers the overall computational complexity of the algorithm and has been used in the experiments reported in Section 4.2.
@@
4 Physical examples
In this final section, we apply the theoretical results and algorithms from Section 3 to two physical systems. We begin with a small system that is small enough to be fully described while still exhibiting all the features of generic systems. We then conclude with a higher-dimensional one.
4.1 A four-mode system
@@
Figure 1: Case $n=4$ and $\text{codim}=3$. Squeezing, $d_{n+j}^{-2}(\omega)$ and anti-squeezing, $d_{i}^{2}(\omega)$ (with $j\in\{1,\ldots,4\}$) spectra, in dB. The specific choice for the parameters gives an avoided crossing between the second ($j=1$) and third ($j=2$) squeezing and anti-squeezing spectra around $\omega=0.794\gamma$ (marked by the vertical black-dashed line) and an avoided crossing between the third ($j=2$) and forth ($j=3$) squeezing spectra around $\omega=0.164\gamma$ (marked by the vertical black-dot-dasjhed line). The inset shows a magnified view of the two avoided crossing regions in the anti-squeezing spectra.
@@
Figure 2: Anti-squeezing spectra in dB ($d_{1}^{2}(\omega,g_{12})$ and $d_{2}^{2}(\omega,g_{12})$) for the case $n=4$, $\mathrm{codim}=3$. (a) For $g_{22}=1.25\gamma$: the surfaces of anti-squeezing avoid crossing through the entire plane $(\omega,g_{11})$. (b) For $g_{22}\approx 1.12\gamma$: the surfaces of squeezing touch at one point of degeneracy in the plane $(\omega,g_{11})$.
@@
Figure 4: Case $n=4$ and $\text{codim}=2$. Squeezing, $d_{n+j}^{-2}(\omega)$ and anti-squeezing, $d_{j}^{2}(\omega)$ (with $j\in\{1,\ldots,4\}$) spectra, in dB. The specific choice for the parameters gives an avoided crossing between the second ($j=1$) and third ($j=2$) squeezing and anti-squeezing spectra around $\omega=0.761\gamma$ (marked by the vertical black-dashed line) and an avoided crossing between the third ($j=2$) and forth ($j=3$) squeezing spectra around $\omega=0.079\gamma$ (marked by the vertical black-dot-dashed line). The inset shows a magnified view of the two avoided crossing regions in the anti-squeezing spectra.
@@
We also considered the situation where the couplings are fully real, replacing $F$ of (29) with its real part. Computing the four curves of singular values, we obtained Figure 4, which is qualitatively similar to Figure 1 and shows two avoided crossings. When investigating the existence of a degeneracy near the second avoided crossing as we did previously, we observed that the Berry phases considered in Theorem 3, and computed via Algorithm 2, took only values that are multiples of $\pi$. This behavior contrasts with the codimension 3 case, where, in general, $\alpha_{j}$ varies smoothly across a continuous range of values when $\theta$ is varied (see Figure 3, particularly $\alpha_{j}$ for $j=2,3$). This suggests that the codimension of degeneracies may be lower than expected  [^5] For instance, the Berry phase takes only values that are multiples of $\pi$ when the matrix function is real symmetric, in which case the codimension of degeneracies is in fact 2, see [13, 12].. Indeed, this turns out to be the case. It is easy to see that, when both $F$ and $G$ are real, the matrix $S(\omega)$ satisfies
@@
Remark 4.
For every value of the parameters $g_{11},g_{22}$ at which the spectrum of the transfer function $S$ displayed a degeneracy, we computed the corresponding spectrum of the Hamiltonian and found no degeneracy. This shows that the spectral degeneracies under consideration in this work manifest themselves only “through the driven-dissipative quantum dynamics”.
@@
4.2 The pulsed microring resonator in synchronously pumped regime
@@
Figure 8: First four most anti-squeezed morphing supermodes ($j={1,\ldots,4}$) for a pulsed microring resonator in the synchronously pumped regime with $n=51$. The enlarged views show the rapid (yet smooth) rotation of two supermodes near an avoided crossing.
@@ 5 Conclusions
Acknowledgements
The authors thank Prof. Luca Dieci for useful discussions. This work has been partially supported by GNCS–INdAM and by the PRIN2022PNRR n. P2022M7JZW “SAFER MESH - Sustainable mAnagement oF watEr Resources ModEls and numerical MetHods" research grant, funded by the Italian Ministry of Universities and Research (MUR), by the European Union through Next Generation EU, M4.C2.1.1, CUP H53D23008930001 and by Plan France 2030 through the project ANR-22-PETQ-0013 (OQuLus).
@@ Appendix A On the joint-MVD of a complex symmetric matrix-valued function depending on a parameter
where $U,V\in\mathbb{C}^{n\times n}$ are unitary and $\Sigma=\operatorname{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{n})$, with non-negative $\sigma_{j}$’s arranged in non-increasing order. If $A^{T}, AA$ also admits a Takagi decomposition
@@ Appendix B Additional figures on rapid variation of the singular vectors
Here appendix we show the behavior of the morphing supermodes of the case studied in subsection 4.1. In particular, Figure 9 consists of eight panels, each corresponding to a fixed value of $m$, and showing the absolute value of the coefficient $U_{m,j}(\omega)$ across all supermodes $j=\{0,\ldots,7\}$. The vertical black dash-dotted line marks the position of the first avoided crossing at $\omega=0.164\gamma$ occurring between the anti-squeezed supermodes $j=2$ and $j=3$. For the symmetry property of the spectrum of the singular values, the same avoided crossing also occurs between the squeezed supermodes $j=6$ and $j=7$. The vertical black dashed line marks the position of the second avoided crossing at $\omega=0.794\gamma$ occurring between the anti-squeezed supermodes $j=1$ and $j=2$. The same avoided crossing also occurs between the squeezed supermodes $j=5$ and $j=6$. Focusing on the top-left panel of the figure, which displays the absolute value of the $m=0$ coefficient for the eight supermodes, we observe a rapid rotation of the anti-squeezed supermode pair $j=2$ (green line) and $j=3$ (red line) at the first avoided crossing, as well as a rotation of the squeezed pair $j=6$ (pink line) and $j=7$ (gray line). At the second avoided crossing, a similar rotation occurs between the anti-squeezed supermodes $j=1$ (orange line) and $j=2$ (green line), and between the squeezed modes $j=5$ (brown line) and $j=6$ (pink line). The same behavior is observed in the panels corresponding to the other coefficients.